Provide all detailed steps to find the limit of the following functions.

Lim x→∞ (e4x - e-2x) ÷ (ln(x+1)


1
Expert's answer
2022-05-03T16:46:00-0400
"\lim\limits_{x\to \infin}\dfrac{e^{4x}-e^{-2x}}{\ln(x+1)}"

Use L'Hôpital's rule


"\lim\limits_{x\to \infin}\dfrac{e^{4x}-e^{-2x}}{\ln(x+1)}=\lim\limits_{x\to \infin}\dfrac{(e^{4x}-e^{-2x})'}{(\ln(x+1))'}"

"=\lim\limits_{x\to \infin}\dfrac{4e^{4x}+2e^{-2x}}{1/(x+1)}=\lim\limits_{x\to \infin}(4e^{4x}+2e^{-2x})(x+1)"

"=\infin"